Law of sines word problem help?
Points A and B are on opposite sides of a river. To find the distance between the points, a third point C is located on the same side of the river as point A. The distance between A and C is 50 feet, ACB is determined to be 47° and BAC is 115°. Find the distance between A and B.
Thank you!

Answer by lwsshak3(11628)

Points A and B are on opposite sides of a river. To find the distance between the points, a third point C is located on the same side of the river as point A. The distance between A and C is 50 feet, ACB is determined to be 47° and BAC is 115°. Find the distance between A and B.
***
Given angle at point C=47˚
Given angle at point A=115˚
So the angle across the river at point B=180-115-47=18˚
Using law of sines:
sin 18/50=sin 47/AB
AB=(sin 47/sin18)*50≈118.34
distance between A and B≈118.34 ft